The displacement of a particle moving on a straight line is given by x=16t-2t^2. Distance travelled by the particle during the first 2 is S1 and during first 6 second is S1. Find 3S2/S1

Grade:11

1 Answers

Vikas TU
6 years ago
At 1 sec:
At 2 sec x = 32-8 = 16m
S1 = 16 m
At 6 sec
S2 = 96 -72 = 24 m
Then 3S2/s1 = 72/16 = 18/4 = 9/2